mmurra00:

Max practices the piano and the guitar for a total of 150 minutes every day. He practices guitar 80 minutes more than he practices the piano. Part A: Write a pair of linear equations to show the relationship between the number of minutes Max practices the guitar (x) and the number of minutes he practices the piano (y) every day. (5 points) Part B: How much time does Max spend practicing the piano every day? (3 points) Part C: Is it possible for Max to have spent 125 minutes practicing the guitar every day? Explain your reasoning. (2 points) Can someone help me through this I don't get it it is for practice?

mxddi3:

so for part a do you have answer choices we can see?

mmurra00:

no we have to find it to find the whole thing ourselves that's whats troubling me

thegoat:

A) x + y = 150 y = x + 80 use substitution x + (x + 80) = 150 2x + 80 = 150 2x = 70 x =35 practicing the guitar then put x back in to solve for y B) y = 35 + 80 = 115 minutes practicing the piano

mmurra00:

thanks

thegoat:

np